# COmoving Computer Acceleration (COCA): $N$-body simulations in an emulated frame of reference

### v1

@@ Abstract
For gravitational $N$-body simulations of dark matter particles, we introduce the COmoving Computer Acceleration (COCA) approach to running cosmological simulations within an emulated frame of reference. While traditional emulators aim to translate initial conditions into final particle positions, directly representing the non-linear dark matter distribution, COCA aims to emulate a frame of reference in which to run a physical simulation with lower computational cost. The approach can be seen as a generalisation and improvement of the idea behind COmoving Lagrangian Acceleration (COLA) [59]. As an illustration, we compare the results of COLA and COCA simulations when forces are evaluated through a particle-mesh (PM) scheme. We find that using our ML-enhanced approach requires very few force evaluations (approximately 8, compared to 20 for COLA) to correct for emulation errors, yielding percent-level accurate power spectra, bispectra, and cross-correlation to a reference simulation.
@@ II.2.   COCA formalism
$\begin{split}\partial_{a}^{2}\boldsymbol{\Psi}_{\mathrm{res}}^{\mathrm{COCA}}(\textbf{q},a)=&-\boldsymbol{\nabla}\Phi(\textbf{x},a)-\partial_{a}^{2}\boldsymbol{\Psi}_{\mathrm{LPT}}(\textbf{q},a)\\ &-\partial_{a}^{2}\boldsymbol{\Psi}_{\mathrm{ML}}(\textbf{q},a).\end{split}$ (7)
@@ III.2.   Scaling of momenta
Figure 3: Slice of the difference between the true momenta of particles p and the LPT prediction $\textbf{p}_{\rm LPT}$, for a test simulation, as a function of scale factor $a$. At late times, the spatial structure of the field $\textbf{p}-\textbf{p}_{\rm LPT}$ remains relatively constant, with most of the time dependence being a simple multiplicative scaling.
In Fig. 3, we plot a slice of the field $\textbf{p}-\textbf{p}_{\rm LPT}$ as a function of scale factor, for one of our test simulations. From visual inspection, we find that the large-scale spatial structure of the field to be emulated does not change significantly as a function of time, particularly at late times, but its magnitude does. We therefore choose to rescale the momenta to be emulated by defining
@@ III.3.   Neural network architecture
For every convolutional layer, we introduce a “style” parameter444 A style parameter in a neural network is an additional input at each layer that encodes dependence on an important feature. In our case, the scale factor $a$ encodes the dependence on clustering at various cosmological times. For more details, we refer the interested reader to Eq. (1)–(3) of Karras et al. [30]. (borrowing the nomenclature from StyleGAN2, 30), where each convolutional kernel is multiplied by an array of the same dimension as the layer’s input and with values equal to the style parameter. Since we are producing a time-dependent emulator, we use the scale factor $a$ as our style parameter. Our network is implemented and trained using a modified version of the map2map [^5] https://github.com/eelregit/map2map/ package and pytorch [50].
@@
We also note that we achieve good performance for our emulator with only $N^{3}=64^{3}$ voxels in the input field, whereas Jamieson et al. [27] uses $N^{3}=128^{3}$ for a similar resolution, resulting in a further factor of 8 reduction in memory requirements. The smaller input size necessitates one fewer resolution layer in our neural network architecture, thus reducing the number of parameters in our model to $2.4\times 10^{6}$, compared to $3.4\times 10^{6}$ in Jamieson et al. [27]. It also requires less padding of the input field: we use periodic padding of 24 voxels, compared to 48 in Jamieson et al. [27].
@@ IV.3.   COCA with misspecified cosmological parameters
Despite the moderate performance of this emulator in the presence of cosmological parameter misspecification, without any force evaluations (i.e., with $n_{\mathrm{f}}=0$), the error on the matter power spectrum would be too large for current cosmological analyses [60]. Therefore, relying solely on an emulator of particles’ trajectories (i.e., a frame of reference emulator with $n_{\mathrm{f}}=0$) as a forward model would produce inappropriate results and would not be a safe use of machine learning. However, trajectories can be rectified in the COCA framework by evaluating gravitational forces and solving for the residual displacements with respect to the emulated frame of reference. In our test, using just 8 force evaluations is sufficient to achieve percent-level agreement in both $P(k)$ and $r^{2}(k)$ for the density field, for all $k<1\,h\,{\rm Mpc}^{-1}$ (see Fig. 11). The same conclusion is true for the velocity field up to $k\approx 0.6\,h\,{\rm Mpc}^{-1}$. Thus, with only a small additional computational cost, we can convert an unsafe use of machine learning in cosmology into a well-behaved one, even when the emulator is applied outside the range of its training data. This is one of the main benefits of COCA compared to traditional emulators of $N$-body simulation results.
@@ IV.4.   COCA versus a Lagrangian displacement field emulator
Figure 12: Relative performance of COCA versus an emulator of the displacement field $\boldsymbol{\Psi}$. We compute the summary statistics outlined in Section III.5 for the final ($a=1$) matter density field and compare results at both the training cosmology and a misspecified one. Although directly emulating $\boldsymbol{\Psi}$ produces a more accurate density field than simply emulating the momentum field p (with $n_{\rm f}=0$), using the COCA framework (emulating the frame of reference and employing additional force evaluations) yields the best performance.
@@ A.1.   Model equations with COCA
With $\textbf{x}(a)=\textbf{x}_{\mathrm{LPT}}(a)+\textbf{x}_{\mathrm{ML}}(a)+\textbf{x}_{\mathrm{res}}(a)$ (denoting the Langrangian perturbation theory, machine-learnt, and residual contributions to the position, respectively), we note for each contribution $\mathrm{y}\in\left\{\mathrm{LPT},\mathrm{ML},\mathrm{res}\right\}$:

### v2

@@ Abstract
For gravitational $N$-body simulations of dark matter particles, we introduce the COmoving Computer Acceleration (COCA) approach to running cosmological simulations within an emulated frame of reference. While traditional emulators aim to translate initial conditions into final particle positions, directly representing the non-linear dark matter distribution, COCA aims to emulate a frame of reference in which to run a physical simulation with lower computational cost. The approach can be seen as a generalisation and improvement of the idea behind COmoving Lagrangian Acceleration (COLA) [59]. As an illustration, we compare the results of COLA and COCA simulations when forces are evaluated through a particle-mesh (PM) scheme. The current implementation can be interpreted either as a means of making COLA more computationally efficient (from a simulator’s perspective) or as a method for correcting emulators of PM-based simulations (from an emulator builder’s perspective). We find that our ML-enhanced approach requires very few force evaluations to correct for emulation errors: it achieves percent-level accurate power spectra, bispectra, and cross-correlation to a reference simulation, with approximately 8 force evaluations, compared to 20 for COLA. Importantly, the COCA formalism is not limited to PM codes. Future iterations of this work will extend the framework to other gravity solvers (e.g., P3M or tree-based methods), with the ultimate goal of improving emulators of these more precise codes.
@@ II.2.   COCA formalism
$\partial_{a}^{2}\boldsymbol{\Psi}_{\mathrm{res}}^{\mathrm{COCA}}(\textbf{q},a)=-\boldsymbol{\nabla}\Phi(\textbf{x},a)-\partial_{a}^{2}\boldsymbol{\Psi}_{\mathrm{LPT}}(\textbf{q},a)-\partial_{a}^{2}\boldsymbol{\Psi}_{\mathrm{ML}}(\textbf{q},a).$ (7)
@@ III.2.   Scaling of momenta
Figure 3: Slice of the difference between the true momenta of particles p and the LPT prediction $\textbf{p}_{\rm LPT}$, for a test simulation, as a function of scale factor $a$. We plot the component orthogonal to the chosen slice. At late times, the spatial structure of the field $\textbf{p}-\textbf{p}_{\rm LPT}$ remains relatively constant, with most of the time dependence being a simple multiplicative scaling.
In Fig. 3, we plot one component of the field $\textbf{p}-\textbf{p}_{\rm LPT}$ as a function of scale factor, for a slice of one of our test simulations. From visual inspection, we find that the large-scale spatial structure of the field to be emulated does not change significantly as a function of time, particularly at late times, but its magnitude does. We therefore choose to rescale the momenta to be emulated by defining
@@ III.3.   Neural network architecture
For every convolutional layer, we introduce a “style” parameter444 A style parameter in a neural network is an additional input at each layer that encodes dependence on an important feature. In our case, the scale factor $a$ encodes the dependence on clustering at various cosmological times. For more details, we refer the interested reader to Eq. (1)–(3) of Karras et al. [30]. (borrowing the nomenclature from StyleGAN2, 30), where each convolutional kernel is multiplied by an array of the same dimension as the layer’s input and with values equal to the style parameter. Since we are producing a time-dependent emulator, we use the scale factor $a$ as a style parameter [similar to 26, who also include a redshift-dependent style parameter]. Our network is implemented and trained using a modified version of the map2map [^5] https://github.com/eelregit/map2map/ package and pytorch [50].
@@
Compared to previous $N$-body emulators, our architecture is most similar to that of Jamieson et al. [27], although we use only $N^{3}=64^{3}$ voxels in the input field, whereas Jamieson et al. [27] employ $N^{3}=128^{3}$ voxels with a similar voxel size. This reduction in voxel count decreases memory requirements by an additional factor of 8 compared to Jamieson et al. [27]. Furthermore, we require less padding of the input field than Jamieson et al. [27]: our approach employs periodic padding of 24 voxels, compared to 48 in their implementation. The smaller input size (resulting from using a scalar field rather than a three-vector field) necessitates one fewer resolution layer in our neural network architecture, thus reducing the number of trainable parameters in our model to $2.4\times 10^{6}$, compared to $3.4\times 10^{6}$ in Jamieson et al. [27]. He et al. [19], who emulate the displacement field of particle-mesh simulations, use $N^{3}=32^{3}$ voxels with the same box size as in our work, resulting in a voxel size that is twice as large. Despite their lower resolution, their architecture employs $8.4\times 10^{6}$ parameters. It is worth noting that while our model has fewer parameters than Jamieson et al. [27], unlike their emulator, our network does not currently encode dependence on cosmological parameters. A similar number of parameters may be necessary once this feature is incorporated.
@@
While it is possible to discuss similarities and differences in the neural network architectures, we cannot directly compare the output of our emulator with that of He et al. [19], Jamieson et al. [27] or Jamieson et al. [26]. This is because these works emulate the Lagrangian displacement field $\boldsymbol{\Psi}$, whereas our emulator predicts the residual momentum $\tilde{\textbf{p}}_{\rm ML}$, the quantity required by the COCA framework. We note that Jamieson et al. [27, 26] employ a more accurate gravity solver in their training simulations compared to the PM model used by He et al. [19] and in this work. However, this difference does not affect the respective neural network architectures.
@@ IV.3.   COCA with misspecified cosmological parameters
Despite the moderate performance of this emulator in the presence of cosmological parameter misspecification, without any force evaluations (i.e., with $n_{\mathrm{f}}=0$), the error on the matter power spectrum would be too large for current cosmological analyses [60]. Therefore, relying solely on an emulator of particles’ trajectories (i.e., a frame of reference emulator with $n_{\mathrm{f}}=0$) as a forward model would produce inappropriate results and would not be a safe use of machine learning. However, trajectories can be rectified in the COCA framework by evaluating gravitational forces and solving for the residual displacements with respect to the emulated frame of reference. In our test, using just 8 force evaluations is sufficient to achieve percent-level agreement in both $P(k)$ and $r^{2}(k)$ for the density field, for all $k<1\,h\,{\rm Mpc}^{-1}$ (see Fig. 11). The same conclusion is true for the velocity field up to $k\approx 0.6\,h\,{\rm Mpc}^{-1}$.
One might question whether an emulator trained on an incorrect cosmology could yield a worse frame of reference than 2LPT at the correct cosmology, thereby negating any advantage of COCA over COLA. We test this question in Fig. 12, where we compare the performance of COCA using a reference frame derived from a misspecified cosmology (while all other computations, including gravitational evolution, employ the correct cosmology as noted above), against COLA, which uses the correct cosmological parameters. For all summary statistics, we find that even with an incorrect cosmology, COCA produces substantially more accurate density fields for any given number of force evaluations. Thus, despite the presence of emulation inaccuracies in the misspecified scenario, our emulator still outperforms the 2LPT frame of reference and significantly reduces the number of force evaluations required to achieve an accurate density field. We have verified that this conclusion remains valid when varying $\Omega_{\rm m}$ within the range $[0.2,0.4]$.
Thus, with only a small additional computational cost, we can convert an unsafe use of machine learning in cosmology into a well-behaved one, even when the emulator is applied outside the range of its training data. This is one of the main benefits of COCA compared to traditional emulators of $N$-body simulation results.
@@ IV.4.   COCA versus a Lagrangian displacement field emulator
Figure 12: Relative performance of COCA and COLA versus an emulator of the displacement field $\boldsymbol{\Psi}$. We compute the summary statistics outlined in Section III.5 for the final ($a=1$) matter density field and compare results at both the training cosmology and a misspecified one for COCA (COLA is evaluated at the correct cosmology in both cases). Although directly emulating $\boldsymbol{\Psi}$ produces a more accurate density field than simply emulating the momentum field p (with $n_{\rm f}=0$), using the COCA framework (emulating the frame of reference and employing additional force evaluations) yields the best performance. Moreover, even when using a misspecified cosmology to predict the frame of reference, COCA significantly outperforms COLA for any given number of time steps.
@@ A.1.   Model equations with COCA
With $\textbf{x}(a)=\textbf{x}_{\mathrm{LPT}}(a)+\textbf{x}_{\mathrm{ML}}(a)+\textbf{x}_{\mathrm{res}}(a)$ (denoting the Lagrangian perturbation theory, machine-learnt, and residual contributions to the position, respectively), we note for each contribution $\mathrm{y}\in\left\{\mathrm{LPT},\mathrm{ML},\mathrm{res}\right\}$: